c. 0.72 moles CHz d. 0.57 moles O2 6. How many grams are represented by: a 6.11 moles CH2O% b. 1.89 moles O2 c. 6.61 moles NaHCO d. 1.14 moles Na
Answer #1

6) mass = (molar mass) x (moles)

a) molar mass of C6H12O6 = (6 x12) + (12 x 1) + ( 6 x 16) = 180 gm/mole

mass = 180 gm/mole x 6.11 mole = 1099.8 gm

b) molar mass of O2 = 2 x 16 = 32 gm/mole

mass = 32 gm/mole x 1.89 mole = 60.48 gm

c) molar mass of NaHCO3 = 23 + 1 + 12 + (3x 16) = 84gm/mole

mass =  84gm/mole x 6.61 mole = 555.24 gm

d) atomic mass of Na = 23 gm/mole

mass = 23 gm/mole x 1.14 mole = 26.22 gm